richie is now 14 y.o., and aimee is 26 y.o. in how many years will the age of richie be 2/3 of aimee's age?: richie is now 14 y.o., and aimee is 26 y.o. in how many years will the age of richie be 2/3 of aimee's age?
Answer by nerdybill(1051)

Let x = yrs when richie is 2/3's aimee's age
.
14+x = (2/3)(26+x)
42+3x = 2(26+x)
42+3x = 52+2x
42+x = 52
x = 10 years